OpenStudy (anonymous):

Rationalize the denominator. Assume variables represent non-negative values.

OpenStudy (anonymous):



OpenStudy (anonymous):

So, we have the cube root of y in the denominator. What would we need to multiply the denominator by to make this rational?

OpenStudy (anonymous):

I really don't know.

OpenStudy (anonymous):

We know: \[(\sqrt[3]{y})^3 = \sqrt[3]{y}\sqrt[3]{y}\sqrt[3]{y} = y\] Using this information, we see that we need to multiply the denominator by the cube root of y, twice.

OpenStudy (anonymous):

Now, if we do the same thing to the numerator, we didn't change the fraction at all. So multiply the top by the same factor.

OpenStudy (anonymous):

Do I multiply it by 2^3?

OpenStudy (anonymous):

Nope. I'll show you this one so you can see what I'm trying to say.

OpenStudy (anonymous):

\[\frac{2}{\sqrt[3]{y}} = \frac{\sqrt[3]{y}\sqrt[3]{y}}{\sqrt[3]{y}\sqrt[3]{y}} \times \frac{2}{\sqrt[3]{y}} = \frac{2 (\sqrt[3]{y})^2}{y}\]

OpenStudy (anonymous):

I got it now.